OpenStudy (anonymous):

find the slope of the line described by x-3y=-6

Nnesha (nnesha):

rewrite that equation into `slope-intercept` form \[\huge\rm y=mx+b\] ^^slope intercept form where m is slope b is y-intercept can you solve that equation for y ?

OpenStudy (anonymous):

3y=x+6

Nnesha (nnesha):

ahh nice! 3 is still at left side :=)

OpenStudy (anonymous):

x+6=3y

Nnesha (nnesha):

hmm same thing how would you cancel out the 3 from right side ?

OpenStudy (anonymous):

subtract x

OpenStudy (anonymous):

from both sides

OpenStudy (phi):

notice you have 3 times y to "undo" the times, think "divide"

OpenStudy (anonymous):

3y divided by 6

Nnesha (nnesha):

you already moved x to the right side \[\huge\rm 3y=x+6\] it's `3 times y` opposite of multiplication is division so divide by 3

OpenStudy (anonymous):

so 3 divided by 6 is 2

Nnesha (nnesha):

we have to solve for y so just divide both sides by 3

Nnesha (nnesha):

\[\huge\rm \frac{ 3y }{ 3 } =\frac{ x+6}{ 3}\]

OpenStudy (anonymous):

x divided by 3 is 0

Nnesha (nnesha):

(x+6)/3 is same as \[\frac{ 1 }{ 3 }x +\frac{ 6 }{ 3 }\]

Nnesha (nnesha):

x divide by 3 is zero ? how ?

Nnesha (nnesha):

there is invisible one \[\huge\rm \frac{ 3y }{ 3 } =\frac{ 1x+6}{ 3}\]

Nnesha (nnesha):

so you are saying 1/3x = 0 nope.

OpenStudy (anonymous):

i thought x was 0 thats why

OpenStudy (anonymous):

i forgot that x equals one

OpenStudy (anonymous):

so the slope is 1/3

Nnesha (nnesha):

no x is just a variable remember we need the equation in slope intercept form \[\huge\rm y=mx+b\]

Nnesha (nnesha):

yes right so what's the y-intercept ?

Nnesha (nnesha):